Grade 12th passMechanics

A cricket ball of mass 150 g moving with a speed of 126km/he hits at the middle of the bat help firmly at its position by the batsman.the ball moves straight back to the bowler after hitting the bat. assuming that collision between ball batsman had to apply to hold the bat firmly at its place would be

Tanu Dey
9 Years agoGrade 12th pass

1 Answer

Ankush
9 Years ago
Here I will assume that collision is elastic there is no loss of kinetic energy. Mass of ball is 150g=0.15kg. Initial speed=126km/h=35m/s. Now as collision is elastic there will be no loss in speed, so velocity will be=-35m/s. Now, initial momentum=mu=0.15×35=5.25. Now final momentum=mv=0.15×(-35)=-5.25. Now foce equals=mv-mu=-5.25-5.25=10.5 therefore batsman had to apply force if magnitude 10.5 newton on the ball in opposite direction of the initial velocity if the ball. He had to apply force oc magnitude 10.5N to hold the vat firmly to its place